OpenStudy (anonymous):

What is the radius of the circle whose circumference is same as that of the area of the rectangle? Here the length is 50 times the breadth of the rectangle of perimeter 1250mm.

OpenStudy (anonymous):

1250=2(b+50b) b=1250/102=12.25 L=50b=612.5 area=612.5*12.25=7 503.125 pir^2=7 503.125 r=sqrt(7 503.125/pi) 2pir=2sqrt(7 503.125*pi)

OpenStudy (anonymous):

@ dipankar - question says numerically, 2*pi*r = area of rectangle hence, the calculated radius should be, radius = 7503.125/(2*pi) = 1194.16

OpenStudy (balrog2000):

Hey all